OpenStudy (anonymous):

simplify 8x+6-2x/x+1

OpenStudy (anonymous):

\[\frac{ 8x+6-2x }{ x+1 }\]

OpenStudy (anonymous):

yah

OpenStudy (ivancsc1996):

\[\frac{ 8x+6-2x }{ x+1 }=\frac{ 6x+6 }{ x+1 }=\frac{ 6(x+1) }{ x+1 }=6\]

OpenStudy (anonymous):

oh cool thanks!

OpenStudy (anonymous):

\[\frac{ 8x -2x +6 }{ x+1 }\] \[\frac{ 6x +6 }{ x+1 }\] \[\frac{ 6( x +1) }{ x+1 }\] \[\frac{ 6 }{ 1 }\] \[6\]